# Tip growth in a strongly concentrated aggregation model follows local geodesics

### v2

@@ 4.2 Grönwall-type bound on the total error
Now we will bound the terms in (4.31) and (4.32) involving $M_{s}$: (A2), (A4) and (B2). First write $I^{j}_{r}-p^{j}_{r}=I^{j}_{\mathbf{c}\left\lfloor r/\mathbf{c}\right\rfloor}-p^{j}_{r}=(I^{j}_{\mathbf{c}\left\lfloor r/\mathbf{c}\right\rfloor}-p^{j}_{\mathbf{c}\left\lfloor r/\mathbf{c}\right\rfloor})+(p^{j}_{\mathbf{c}\left\lfloor r/\mathbf{c}\right\rfloor}-p^{j}_{r})$. Let $M_{s}^{\prime}=\int_{s}^{t}(I^{j}_{\mathbf{c}\left\lfloor r/\mathbf{c}\right\rfloor}-p^{j}_{\mathbf{c}\left\lfloor r/\mathbf{c}\right\rfloor})C_{r}\,\mathrm{d}r$, then by Lemma 4.4 we have $|p^{j}_{\mathbf{c}\left\lfloor r/\mathbf{c}\right\rfloor}-p^{j}_{r}|\leq A\mathbf{c}$ for all $r\in[0,T]$, so
@@
We will choose a particular sequence $(s_{m})_{1\leq m\leq N_{t}}$ in $[0,t-10\mathbf{c}]$ such that $\bigcap_{m=1}^{N_{t}}E_{s_{m}}$ holds with high probability, and on this event $|M_{s}^{\prime}|$ is bounded uniformly on $[0,t-10\mathbf{c}]$. Let $r_{1}=10\mathbf{c}$ and $s_{1}=t-r_{1}$. for $m\geq 1$ set $r_{m+1}=r_{m}+\mathbf{c}^{\frac{1}{R}}\sqrt{r_{m}}$ and $s_{m}=\max(t-r_{m},0)$. Then $r_{m+1}-r_{m}\geq\mathbf{c}^{\frac{1}{R}}\sqrt{r_{m}-r_{m-1}}$, and so $r_{m+1}-r_{m}\geq\mathbf{c}^{\frac{1}{2^{m}}+\frac{1}{R}(2-\frac{1}{2^{m}})}\geq\mathbf{c}^{\frac{4}{R}}$ for $m$ greater than some constant $m_{R}$. Hence $N_{t}:=\min\{m:r_{m}\geq t\}\leq A\mathbf{c}^{-\frac{4}{R}}$. Let $s\in(s_{m+1},s_{m})$ for some $m\geq 1$, then
@@
Propositions 4.7 and 4.8 give us $\delta_{\mathrm{total}}(t)\leq A\int_{0}^{t}\frac{\delta_{\mathrm{total}}(s)}{(t-s)^{51/100}}\,\mathrm{d}s+A\mathbf{c}^{1/2R}$ for all $t\in[0,T]$. Using a singular version of Grönwall’s inequality (see [15, Corollary 2]), this implies $\delta_{\mathrm{total}}(T)\leq A\mathbf{c}^{1/2R}$ as required. ∎
@@ 5 Proof of main theorem
Acknowledgements
The initial version of this paper was completed while studying for a PhD at Lancaster University, and many thanks are due to my then-supervisor Amanda Turner for her guidance and patience. Thanks also to both examiners Dmitry Belyaev and Dmitry Korshunov who provided useful feedback. I benefited significantly from discussions with Alan Sola and Fredrik Viklund while attending the programme “The Analysis and Geometry of Random Spaces” at MSRI.

### v3

@@ 5.2 Grönwall-type bound on the total error
Now we will bound the terms in (5.28) and (5.29) involving $M_{s}$: (A2), (A4) and (B2). First write $I^{j}_{r}-p^{j}_{r}=I^{j}_{\mathbf{c}\left\lfloor r/\mathbf{c}\right\rfloor}-p^{j}_{r}=(I^{j}_{\mathbf{c}\left\lfloor r/\mathbf{c}\right\rfloor}-p^{j}_{\mathbf{c}\left\lfloor r/\mathbf{c}\right\rfloor})+(p^{j}_{\mathbf{c}\left\lfloor r/\mathbf{c}\right\rfloor}-p^{j}_{r})$. Let
$M_{s}^{\prime}:=\int_{s}^{t}(I^{j}_{\mathbf{c}\left\lfloor r/\mathbf{c}\right\rfloor}-p^{j}_{\mathbf{c}\left\lfloor r/\mathbf{c}\right\rfloor})C_{r}\,\mathrm{d}r,$
then by Proposition 5.4 we have $|p^{j}_{\mathbf{c}\left\lfloor r/\mathbf{c}\right\rfloor}-p^{j}_{r}|\leq A\mathbf{c}$ for all $r\in[0,T]$, so
@@
We will choose a particular sequence $(s_{m})_{1\leq m\leq N_{t}}$ in $0,t-10\mathbf{c}]$ such that $\bigcap_{m=1}^{N_{t}}E_{s_{m}}$ holds with high probability, and on this event $|M_{s}^{\prime}|$ is bounded uniformly on $0,t-10\mathbf{c}]$. Let $r_{1}=10\mathbf{c}$ and $s_{1}=t-r_{1}$. for $m\geq 1$ set $r_{m+1}=r_{m}+\mathbf{c}^{\frac{1}{R}}\sqrt{r_{m}}$ and $s_{m}=\max(t-r_{m},0)$. Then $r_{m+1}-r_{m}\geq\mathbf{c}^{\frac{1}{R}}\sqrt{r_{m}-r_{m-1}}$, and so $r_{m+1}-r_{m}\geq\mathbf{c}^{\frac{1}{2^{m}}+\frac{1}{R}(2-\frac{1}{2^{m}})}\geq\mathbf{c}^{\frac{4}{R}}$ for $m$ greater than some constant $m_{R}$. Let $N_{t}:=\min\{m:r_{m}\geq t\}$ and it follows from the above that $N_{t}\leq A\mathbf{c}^{-\frac{4}{R}}$. Let $s\in(s_{m+1},s_{m})$ for some $m\geq 1$, then
@@
Propositions 5.5 and 5.6 give us $\delta_{\mathrm{total}}(t)\leq A\int_{0}^{t}\frac{\delta_{\mathrm{total}}(s)}{(t-s)^{51/100}}\,\mathrm{d}s+A\mathbf{c}^{1/2R}$ for all $t\in[0,T]$. Using a singular version of Grönwall’s inequality (see [16, Corollary 2]), this implies $\delta_{\mathrm{total}}(T)\leq A\mathbf{c}^{1/2R}$ as required. ∎
@@ 6 Proof of main theorem
Since $\mathbf{c}^{-1}D=o(1)$, this upper bound tends to zero in probability as $\mathbf{c}\to 0$, and $\mathbb{P}(E_{1}\cap E_{2})\to 1$, giving us $d_{\mathrm{BW}}(\delta_{\xi_{t}}\otimes m_{[0,T]},\bar{\mu}_{t}\otimes m_{[0,T]})\overset{p}{\to}0$ as required. ∎
Acknowledgements
The initial version of this paper was completed while studying for a PhD at Lancaster University, and many thanks are due to my then-supervisor Amanda Turner for her guidance and patience. Thanks also to both examiners Dmitry Belyaev and Dmitry Korshunov who provided useful feedback. I benefited significantly from discussions with Alan Sola and Fredrik Viklund. An anonymous reviewer gave many suggestions which improved the presentation.
Declarations